# Targeting Alignment: Extracting Safety Classifiers of Aligned LLMs

### v2

@@ Abstract
Alignment in large language models (LLMs) is used to enforce guidelines such as safety. Yet, alignment fails in the face of jailbreak attacks that modify inputs to induce unsafe outputs. In this paper, we introduce and evaluate a new technique for jailbreak attacks. We observe that alignment embeds a safety classifier in the LLM responsible for deciding between refusal and compliance, and seek to extract an approximation of this classifier: a surrogate classifier. To this end, we build candidate classifiers from subsets of the LLM. We first evaluate the degree to which candidate classifiers approximate the LLM’s safety classifier in benign and adversarial settings. Then, we attack the candidates and measure how well the resulting adversarial inputs transfer to the LLM. Our evaluation shows that the best candidates achieve accurate agreement (an $F_{1}$ score above 80%) using as little as 20% of the model architecture. Further, we find that attacks mounted on the surrogate classifiers can be transferred to the LLM with high success. For example, a surrogate using only 50% of the Llama 2 model achieved an attack success rate (ASR) of 70% with half the memory footprint and runtime—a substantial improvement over attacking the LLM directly, where we only observed a 22% ASR. These results show that extracting surrogate classifiers is an effective and efficient means for modeling (and therein addressing) the vulnerability of aligned models to jailbreaking attacks.
1 Introduction
@@
Improvements in machine learning—specifically the introduction of the transformer architectures [47]—have led to increasingly powerful large language models (LLMs). Whether they are accessed through an API or interface (e.g., GPT [33], Claude [1]) or published as open-source (e.g., Llama 2 [30], Llama 3 [31], Qwen [46], etc.), these models have become the de facto tool for tasks involving natural language. They serve as foundations for new tools, in which they are augmented with new capabilities [2, 32] or fine-tuned on a downstream task [16]. In adapting these models to tasks, LLMs undergo an alignment process in which they are further trained (refined) to satisfy safety objectives [30, 39]. For example, alignment guidelines often prevent a model from outputting responses that would be offensive, discriminatory, or harmful (i.e., unsafe). It is acknowledged that alignment can (and often does) fail in adversarial settings, resulting in a jailbreak: an unsafe input being accepted as safe. Indeed, many studies have explored the automated generation of jailbreak inputs, i.e., adversarial examples [65, 27, 25]. However, these attacks, specifically in white-box settings, suffer from two major drawbacks: they are inefficient (i.e., high runtime and memory footprint) and employ various heuristics (e.g., maximize the likelihood of a specific output sentence [65], apply fixed changes to the input [53, 20], etc.) that limit the search space of adversarial inputs and thus lower their efficacy. Thus, these flaws limit the potential of attacks to assess the robustness of models.
@@
We perform a detailed analysis of the efficacy of our approach over 4 models and 2 datasets. We first evaluate candidates with respect to benign inputs, assessing whether each candidate classifier agrees with the safety classification of the LLM model. We find that the best candidates achieve accurate agreement (an $F_{1}$ score above 80%) using as little as 20% of the model architecture. Next, we evaluate the candidate classifiers with respect to transferability of adversarial inputs (i.e., as generated using best-in-class jailbreaking algorithms) from the candidates onto the LLM. These latter experiments show that some candidate classifiers are extremely effective and efficient tools for targeting LLMs. For example, a surrogate classifier using only 50% of the Llama 2 model achieved an attack success rate (ASR) of 70%—a substantial improvement over attacking the LLM directly, where we observed a 22% ASR—with less than half the runtime (350s/sample against 850s/sample) and the memory footprint (14 GB of VRAM usage against 31 GB).
@@
2 Background
2.1 Large Language Models (LLMs)
@@ 2.2 Alignment and Safety
Alignment. Since the training data used for LLMs comes from diverse sources, biases and undesirable behaviors emerge [14], leading to misaligned pretrained models [3]. In order to prevent those behaviors from happening, models often go through an alignment process that regulates their outputs according to given guidelines. Several methods exist and are not mutually exclusive [30]. Approaches such as supervised fine-tuning (SFT) use human-generated sample responses, while reinforcement learning with human feedback (RLHF) [4, 8] trains a neural network that acts as a reward for the LLM. Since training another reward model may be costly, direct preference optimization (DPO) [39] uses the LLM as its own reward model.
@@
Guidelines and taxonomies. The existence of alignment as a technique is a direct product of the complexity of identifying what is precisely expected from models (i.e., their objective). Safety is a broad term; therefore, previous work has identified taxonomies [18] that represent a wide range of unsafe behaviors. This allows researchers to evaluate alignment on a finer scale and understand where it could be improved. For example, when the prompts are decomposed into different categories such as self-harm, privacy, harassment, etc., models tend to refuse the prompts related to self-harm and accept the prompt related to privacy [10].
@@
Identifying unsafe outputs. Classifying LLM outputs is challenging. Since aligned models tend to create strong refusal responses for detected unsafe prompts, a naive solution is to test whether certain refusal keywords are in the output (e.g., ”Sorry”, ”as a responsible AI”, etc.) [65]. Alternatively, judge models (also LLMs) are trained to classify the safety of outputs with more [18] or less [29] granularity on the classes they predict, similarly to the taxonomies explained above. A limitation of judge models is that they can also be attacked [28, 40], leading to a possible high rate of false negatives (unsafe outputs judged as safe).
2.3 Attacks on LLMs
@@
LLMs, like any machine learning model, are prone to attacks. While several exploits exist (e.g., model-stealing [5], text misclassification, etc. [61]), the jailbreak exploit remains the most prevalent. Jailbreak refers to the compliance of an aligned LLM for an unsafe query, as defined by the alignment guidelines. Jailbreak attacks fall into one of two threat models: the white-box threat model, in which model parameters are known, and the black-box threat model, in which only access to the output of the model is assumed.
@@
White-box. These attacks have been studied on open-weights models such has Llama 2 [30], Llama 3 [31], or Vicuna [7]. The most prevalent, GCG [65], uses greedy coordinate gradient (GCG) to mitigate the absence of mapping from embeddings (i.e., vectors) to tokens. Following this, many new white-box algorithms have been introduced [52, 9], improving existing algorithms [55, 37, 25, 26], using different objectives [57], or using specific properties of the model (e.g., logits [24], special tokens [58], or generation [17]). Attacks like GCG have an average attack success rate (ASR) above 50% [29] on most open-source models.
@@
Black-box. These attacks have been widely studied, given their prevalence in practical scenarios [33, 1]. Most of them are driven by empirical heuristics: they apply transformations to the initial prompt such as ASCII-art [20], translation into underrepresented languages [53], or using pre-existing templates [54]. Black-box attacks achieve remarkable efficacy. For instance, ArtPrompt [20] can achieve ASR as high as 80% on commercial LLMs such has GPT-3.5.
3 Methodology
@@ 3.1 Preliminaries
Motivation of existence. Refusal from aligned models often comes with a set of predefined outputs (e.g., “I cannot fulfill your request. I’m just an AI”) as a product of alignment. As the space of refusal outputs is very limited, such outputs can be easily detected. We assume the following: the first generated token after the input is enough information to know whether the output will be a refusal or not. Under this assumption, there exists a set of refusal tokens $V_{r}$ such that, for a given input sequence of tokens $x=x_{1}x_{2}\dots x_{T}\in V^{\ast}$, $x_{T+1}\in V_{r}$ if and only if the output answer is a refusal. Thus, classifying unsafe inputs reduces to verifying $x_{T+1}\in V_{r}$. In a deterministic generation (i.e., without sampling), this is done by finding the token $x_{T+1}$ that satisfies $x_{T+1}=\arg\max\limits_{i\in V}\text{softmax}(Ah^{(D)}_{T})_{i}=\arg\max\limits_{i\in V}[Ah^{(D)}_{T}]_{i}$. As explained in Section 2.1, the last embedding (after the last decoder) $h_{T}^{(D)}$ entirely determines the next token. Moreover, the sequence of decoders in the model transforms the sequence of embeddings such that for a given decoder $i$, the $t$-th element $h_{t}^{(i)}$ is a function of $h_{1}^{(i-1)},\dots,h_{t}^{(i-1)}$. Thus, the partial information on the classification contained in $h_{1}^{(i-1)},\dots,h_{T}^{(i-1)}$ propagates to $h_{T}^{(i)}$ until it reaches $h_{T}^{(D)}$.
@@
This score ranges from 1-1 to $1$ where scores below 0 indicate that representations overlap, and scores above 0.25 or 0.5 imply a weak or reasonable separation [41].
@@
Result. We measure this score on four LLMs and two datasets (see Section 4.1 for the detailed experimental setup). Figure 2 shows the silhouette score of the embeddings for different models and datasets, with respect to the normalized decoder position (ratio between the current decoder position and the total number of decoders).
While it would be expected to see a nondecreasing trend [^1] Usually, the subsequent layers of classifiers do not lose information on the classification, hence a nondecreasing trend. For LLMs, this classification is only one of many tasks that are performed, which is why we observe a compression of the linear information in later layers., we see that after the silhouette score reaches its maximum, it starts to decrease. A possible explanation for this phenomenon is rooted in the fact that models are designed to perform multiple tasks. Therefore, after extracting and integrating the safety information from the input, the other tasks add more information to the embedding that is not related to safety. Thus, the safety separation decreases (i.e., the linear information is compressed), but remains high enough for the classification. Besides the difference in magnitude of separation between the two datasets, we note that the trend is exactly the same for all models.
@@
3.2 Threat Model
Motivated by our preliminary results, our adversary aims to extract an approximation of the safety classifier of a target aligned model, called a surrogate classifier. A successful extraction would allow subsequent attacks to be made with more effective white-box jailbreak approaches, while forgoing the complex computational costs they incur, as we discuss in Section 5.2. We will extract the surrogate classifier using subsets of the model, thus white-box access is assumed (i.e., access to the model weights and architecture).
@@ 3.2 Threat Model
Figure 3: Methodology overview. In the first step, we extract the safety classifier of an LLM by (A) selecting a structure within the model and (B) training a classification head on the predictions of the LLM to create a candidate classifier. We verify its performance in benign setting. Then, we measure the transferability between the LLM and the candidate in both ways, i.e., adversarial examples of the LLM (C) transfer to the candidate classifiers and adversarial examples of the candidate classifiers (D) transfer to the LLM.
3.3 Problem Formulation
@@
$\mathcal{R}(f_{\theta}(x))=\mathcal{C}(f_{\theta^{\ast}}(x))$ (3)
@@ 3.4 Extracting the Classifier
    Collect data points from the structure $(f_{\theta^{\ast}}(x),\mathcal{R}(f_{\theta}(x)))$ consisting in the features extracted from the structure and the predicted label of the LLM.
@@
Structures. As explained in Section 2.1, most LLMs are built as sequences of decoders: processing units that take a sequence of vectors and output another sequence of vectors of the same length. LLMs can be scaled either in depth by adding decoders or in width by increasing the dimensionality of embeddings. Therefore, there are many ways to “cut” the model and thus many possible structures. It is clear that there is a separation between unsafe and safe representations at the end of aligned models [26] largely due to the strong refusal outputs. In addition, previous work [23, 59] has shown that separation occurs in early decoders [23]. Thus, we estimate the location of the safety classifier via two hypotheses. First, the classifier is at the decoder level. Second, the information relevant to the classification at a given decoder is contained at the end of the sequence (as discussed in Section 3.1).
@@
By construction, the input space of $f_{i,i+\delta}$ depends on the output space of $f_{1,i}$, since the model was trained with the entire architecture. In this work, we set $i$ to 1. We leave the exploration of structures that isolate intermediate decoders to future work (see Section 5.3) because they would require techniques to mitigate the absence of earlier decoders (leading to a lower accuracy of the extraction), which is out of the scope of this work.
Finally, $\delta$ will be termed the candidate size and we will refer to $\frac{\delta}{D}$ as the normalized candidate size (also referred to as percentage of the corresponding LLM), i.e., the proportion of decoders of the LLM used by the candidate.
Training. For the selected structure $f_{i,i+\delta}$, the goal is to train a classification head $\mathcal{C}$ to obtain a candidate classifier. To avoid overfitting due to a complex classification head, our approach leverages linear probing, which has shown promising results for safety representation [21, 56]. The first step is to create a dataset $\mathcal{X}=(f_{i,i+\delta}(x),\mathcal{R}(f_{\theta}(x)))$ consisting in the extracted feature from the structure $f_{i,i+\delta}$ and the predicted label [^2] Since we focus on extracting the classifier and not improve it, it is beneficial to have both right and wrong labels from the classifier. by applying the classification rule $\mathcal{R}$ to the output of the LLM $f_{\theta}$ for each prompt $x$. Finally, the classification head is trained with this objective:
@@
Once the classification head is trained, we obtain a candidate classifier $\mathcal{C}\circ f_{i,i+\delta}$ of size $\delta$.
3.5 Efficient Jailbreaking
As a reminder, our goal is to induce a jailbreak in the LLM using candidate classifiers as approximations of the embedded safety classifier. This involves two steps: generating adversarial samples from the candidate classifiers and transferring them to the LLM. Although similar to transferability-based attacks that use source models different from the target model to craft adversarial examples [34, 35], our approach uses a subset of the target model (the LLM) as the source, using the insights from Section 3.1 as foundations for a higher efficiency (and efficacy) of the attack.
We attack the candidate classifiers using the well established GCG attack [65]. We detail below how we modify this attack to make it usable for a classifier.
From heuristics to misclassification. The original adversarial objective of GCG is to maximize the likelihood of a given target sequence (e.g., “Sure, here is how to build a bomb.”) for a given input $x$ on which $A$ tokens are added as an adversarial suffix, i.e.,

### v3

@@ Abstract
Alignment in large language models (LLMs) is used to enforce guidelines such as safety. Yet, alignment fails in the face of jailbreak attacks that modify inputs to induce unsafe outputs. In this paper, we introduce and evaluate a new technique for jailbreak attacks. We observe that alignment embeds a safety classifier in the LLM responsible for deciding between refusal and compliance, and seek to extract an approximation of this classifier: a surrogate classifier. To this end, we build candidate classifiers from subsets of the LLM. We first evaluate the degree to which candidate classifiers approximate the LLM’s safety classifier in benign and adversarial settings. Then, we attack the candidates and measure how well the resulting adversarial inputs transfer to the LLM. Our evaluation shows that the best candidates achieve accurate agreement (an $F_{1}$ score above 80%) using as little as 20% of the model architecture. Further, we find that attacks mounted on the surrogate classifiers can be transferred to the LLM with high success. For example, a surrogate using only 50% of the Llama 2 model achieved an attack success rate (ASR) of 70% with half the memory footprint and runtime—a substantial improvement over attacking the LLM directly, where we only observed a 22% ASR. These results show that extracting surrogate classifiers is an effective and efficient means for modeling (and therein addressing) the vulnerability of aligned models to jailbreaking attacks. The code is available at https://github.com/jcnf0/targeting-alignment.
I Introduction
@@
Improvements in machine learning—specifically the introduction of the transformer architectures [50]—have led to increasingly powerful large language models (LLMs). Whether they are accessed through an API or interface (e.g., GPT [35], Claude [1]) or published as open-source (e.g., Llama 2 [32], Llama 3 [33], Qwen [49], etc.), these models have become the de facto tool for tasks involving natural language. They serve as foundations for new tools, in which they are augmented with new capabilities [2, 34] or fine-tuned on a downstream task [17]. In adapting these models to tasks, LLMs undergo an alignment process in which they are further trained (refined) to satisfy safety objectives [32, 42]. For example, alignment guidelines often prevent a model from generating responses that would be offensive, discriminatory, or harmful (i.e., unsafe). It is acknowledged that alignment can (and often does) fail in adversarial settings, resulting in a jailbreak: an unsafe input being accepted as safe. Indeed, many studies have explored the automated generation of jailbreak inputs, i.e., adversarial examples [70, 28, 26]. However, these attacks, specifically in white-box settings, suffer from two major drawbacks: they are inefficient (i.e., high runtime and memory footprint) and employ various heuristics (e.g., maximize the likelihood of a specific output sentence [70], apply fixed changes to the input [57, 21], etc.) that limit the search space of adversarial inputs and thus lower their efficacy and their ability to assess the robustness of models.
@@
We perform a detailed analysis of the efficacy of our approach over 4 models and 2 datasets. We first evaluate candidates with respect to benign inputs, assessing whether each candidate classifier agrees with the safety classification of the LLM. We find that the best candidates achieve accurate agreement (an $F_{1}$ score above 80%) using as little as 20% of the model architecture. Next, we evaluate the candidate classifiers with respect to transferability of adversarial inputs (i.e., as generated using best-in-class jailbreaking algorithms) from the candidates onto the LLM. These latter experiments show that some candidate classifiers are extremely effective and efficient tools for targeting LLMs. For example, a surrogate classifier using only 50% of the Llama 2 model achieved an attack success rate (ASR) of 70%—a substantial improvement over attacking the LLM directly, where we observed a 22% ASR—with less than half the runtime (350s/sample against 850s/sample) and the memory footprint (14 GB of VRAM usage against 31 GB).
@@
II Background
II-A Large Language Models (LLMs)
@@ II-B Alignment and Safety
Alignment. Since the training data used for LLMs comes from diverse sources, biases and undesirable behaviors emerge [15], leading to misaligned pretrained models [3]. In order to prevent those behaviors from happening, models often go through an alignment process that regulates their outputs according to given guidelines. Several methods exist and are not mutually exclusive [32]. Approaches such as supervised fine-tuning (SFT) use human-generated sample responses, while reinforcement learning with human feedback (RLHF) [4, 9] trains a neural network that acts as a reward for the LLM. Since training another reward model may be costly, other approaches such as direct preference optimization (DPO) [42] uses the LLM as its own reward model.
@@
Guidelines and taxonomies. The existence of alignment as a technique is a direct product of the complexity of identifying what is precisely expected from models (i.e., their objective). Safety is a broad term; therefore, previous work has identified taxonomies [19] that represent a wide range of unsafe behaviors. This allows researchers to evaluate alignment on a finer scale and understand where it could be improved. For example, when the prompts are decomposed into different categories such as self-harm, privacy, or harassment, models tend to refuse prompts related to self-harm and accept prompts related to privacy [11].
@@
Identifying unsafe outputs. Classifying LLM outputs is challenging. Since earlier aligned models created strong refusal responses, a naive solution is to test whether certain refusal keywords are in the output (e.g., ”Sorry”, ”as a responsible AI”, etc.) [70]. However, this method is less accurate for more recent models that were aligned with a different approach than refusal training [59]. Alternatively, LLM judge models are trained to classify the safety of outputs with more [19] or less [30] granularity on the classes they predict, similarly to the taxonomies explained above. A limitation of judge models is that they can also be attacked [29, 43], leading to a possible high rate of false negatives (unsafe outputs judged as safe).
II-C Attacks on LLMs
@@
LLMs, like any machine learning model, are prone to attacks. While several exploits exist (e.g., model-stealing [5], text misclassification, etc. [66]), the jailbreak exploit remains the most prevalent. Jailbreak refers to the compliance of an aligned LLM for an unsafe query, as defined by the alignment guidelines. Jailbreak attacks fall into one of two threat models: the white-box threat model, in which model parameters are known, and the black-box threat model, in which only access to the output of the model is assumed. Beyond their use for malicious behavior, jailbreak attacks are also used to test LLMs before their deployment through red-teaming [51].
@@
White-box. These attacks have been studied on open-weights models such as Llama 2 [32], Llama 3 [33], or Vicuna [8]. One of the first white-box attacks, GCG [70], uses greedy coordinate gradient (GCG) to mitigate the absence of mapping from embeddings (i.e., vectors) to tokens. Following this, many new white-box algorithms have been introduced [56, 10], improving existing algorithms [60, 39, 26, 27], using different objectives [62], or using specific properties of the model (e.g., logits [25], special tokens [63], or generation [18]). Attacks like GCG have an average attack success rate (ASR) above 50% [30] on most open-source models.
@@
Black-box. These attacks have been widely studied, given their prevalence in practical scenarios [35, 1]. These attacks generally employ heuristics (transformations to the initial prompt such as ASCII-art [21], translation into underrepresented languages [57], or using pre-existing templates [58]), search algorithms [7, 31], or a combination thereof. Black-box attacks achieve remarkable efficacy. For instance, ArtPrompt [21] can achieve ASR as high as 80% on commercial LLMs such as GPT-3.5.
III Methodology
@@ III-A Empirical Evidence of a Safety Classifier
Motivation of existence. While refusal from aligned models started as sets of predefined outputs (e.g., “I cannot fulfill your request. I’m just an AI”), this idea was challenged [41] and recent efforts [59] have tried to pivot away from refusal training and focus on safe-completions. Regardless, given an input, the model has to decide whether an input is safe or unsafe. We assume the following: the last intermediate state before the output is enough information to know whether the output will be a refusal or not. Under this assumption, there exists a set of refusal intermediate states $H_{r}$ such that, for a given input sequence of tokens $x=x_{1}x_{2}\dots x_{T}\in V^{\ast}$, $h^{(D)}_{T}\in H_{r}$ if and only if the output answer is a refusal. Thus, classifying unsafe inputs reduces to verifying $h^{(D)}_{T}\in H_{r}$.
@@
Silhouette scores range from 1-1 to $1$ where scores below 0 indicate that representations overlap, and scores above 0.25 or 0.5 imply a weak or reasonable separation [44].
@@
Result. We measure this score on four LLMs and two datasets (see Section IV-A for the detailed experimental setup). Figure 2 shows the silhouette score of the embeddings for different models and datasets, with respect to the normalized decoder position (ratio between the current decoder position and the total number of decoders).
While it would be expected to see a nondecreasing trend [^1] Usually, the subsequent layers of classifiers do not lose information on the classification, hence a nondecreasing trend. For LLMs, this classification is only one of many tasks that are performed, which is why we observe a compression of the linear information in later layers., we see that after the silhouette score reaches its maximum, it starts to decrease. A possible explanation for this phenomenon is rooted in the fact that models are designed to perform multiple tasks. Therefore, after extracting and integrating the safety information from the input, the other tasks add more information to the embedding that is not related to safety. Thus, the safety separation decreases (i.e., the linear information is compressed), but remains high enough for the classification. Besides the difference in magnitude of separation between the two datasets, the trend is the same for all models.
@@
III-B Threat Model
Motivated by our preliminary results, our adversary aims to extract an approximation of the safety classifier of a target aligned model, called a surrogate classifier. A successful extraction would allow subsequent attacks to be made with more effective white-box jailbreak approaches, while forgoing the complex computational costs they incur, as we discuss in Section V-B. We will extract the surrogate classifier using subsets of the model, thus white-box access is assumed (i.e., access to the model weights). In practice, the actor could be a red-teamer that would benefit from the efficiency and efficacy gains of the approach when auditing a model.
@@ III-B Threat Model
Figure 3: Methodology overview. In the first step, we extract the safety classifier of an LLM by (A) selecting a structure within the model and (B) training a classification head on the predictions of the LLM to create a candidate classifier. We verify its performance in benign settings. Then, we measure the transferability between the LLM and the candidate in both ways: adversarial examples of the LLM to the candidate (C) and adversarial examples of the candidate to the LLM (D).
III-C Problem Formulation
@@
$\mathcal{R}(f_{\theta}(x))=\mathcal{C}(f_{\theta^{\ast}}(x)).$ (3)
@@ III-D Extracting the Classifier
    Collect data points from the structure $(f_{\theta^{\ast}}(x),\mathcal{R}(f_{\theta}(x)))$ consisting of the features extracted from the structure and the predicted label of the LLM.
@@
Structures. As explained in Section II-A, most LLMs are built as sequences of decoders: processing units that take a sequence of vectors and output another sequence of vectors of the same length. LLMs can be scaled either in depth by adding decoders or in width by increasing the dimensionality of embeddings. Therefore, there are many ways to “cut” the model and thus many possible structures. It is clear that there is a separation between unsafe and safe representations at the end of aligned models [27] largely due to the strong refusal outputs. In addition, previous work [24, 64] has shown that separation occurs in early decoders [24]. Thus, we estimate the location of the safety classifier via two hypotheses. First, the classifier is at the decoder level. Second, the information relevant to the classification at a given decoder is contained at the end of the sequence (as discussed in Section III-A).
@@
By construction, the input space of $f_{i,i+\delta}$ depends on the output space of $f_{1,i}$, since the model was trained with the entire architecture. In this work, we set $i$ to 1. We leave the exploration of structures that isolate intermediate decoders to future work (see Section V-C) because they would require techniques to mitigate the absence of earlier decoders, which is out of the scope of this work.
Finally, $\delta$ will be termed the candidate size and we will refer to $\frac{\delta}{D}$ as the normalized candidate size (also referred to as percentage of the corresponding LLM), i.e., the proportion of decoders of the LLM used by the candidate. We will focus on the latter when comparing different LLMs as the number of parameters is spread evenly across decoders.
Training. For the selected structure $f_{i,i+\delta}$, the goal is to train a classification head $\mathcal{C}$ to obtain a candidate classifier. To avoid overfitting due to a complex classification head, our approach leverages linear probing, which has shown promising results for safety representation [22, 61]. The first step is to create a dataset $\mathcal{X}=(f_{i,i+\delta}(x),\mathcal{R}(f_{\theta}(x)))$ consisting of the extracted feature from the structure $f_{i,i+\delta}$ and the predicted label [^2] Since we focus on extracting the classifier and not improve it, it is beneficial to have both right and wrong labels from the classifier. by applying the classification rule $\mathcal{R}$ to the output of the LLM $f_{\theta}$ for each prompt $x$. Finally, the classification head is trained using the following objective:
@@
Once the classification head is trained, we obtain a candidate classifier noted $\mathcal{C}\circ f_{i,i+\delta}$ of candidate size $\delta$.
III-E Efficient Jailbreaking
As a reminder, our goal is to induce a jailbreak in the LLM using candidate classifiers as approximations of the embedded safety classifier. This involves two steps: generating adversarial samples from the candidate classifiers and transferring them to the LLM. Although similar to transferability-based attacks that use source models different from the target model to craft adversarial examples [37, 36], our approach uses a subset of the target model (the LLM) as the source, using the insights from Section III-A as foundations for a higher efficiency (and efficacy) of the attack.
We attack the candidate classifiers using the well established GCG attack [70]. Since the attack normally targets an LLM that outputs a sequence of tokens, some adjustments to its objective are required to make it usable on a candidate classifier, which we detail below.
From heuristics to misclassification. The original adversarial objective of GCG is to maximize the likelihood of a target sequence (e.g., “Sure, here is how to build a bomb.”) for an input $x$ on which $A$ tokens are added as an adversarial suffix: